In a town election, the winning candidate had 220 more votes than the loser. If 810 votes were castmin all, how many votes did each candidate receive?
Answer by ptaylor(2198)

Let x=number of votes for the loser
Then 220+x=number of votes for the winner
Loser votes+winner votes=total votes, So:
x+220+x=810 subtract 220 from both sides
x+220-220+x=810-220 collect like terms
2x=590 divide both sides by 2
x=295---------------------------number of votes for loser
220+x=220+295=515---------------number of votes for winner

Ck
295+515=810
810=810
Hope this helps----ptaylor